OpenStudy (anonymous):

what equation is equivalent to 2/3(9x-3)=5(2x+7)

OpenStudy (anonymous):

\[2(9x-3)=15(2x+7)\\18x-6=30x+105\\-12x=111\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problemo